Yanmei has contributed $250 to an RRSP at the end of each 3-month period for the past 35 years. During this time, the RRSP has earned an average of 11.5%/a compounded quarterly.

a) How much will the RRSP be worth at maturity?

b) How much of the investment will be interest earned over the 35 years?


1
Expert's answer
2021-02-24T06:42:32-0500

"Solution:~ Given ~ that ~ PMT=\$250 
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~n=4 ~ for ~ querterly ~ payment 
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~t=35~ years
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~r=11.5\% =0.115
\\a) We~have ~ to ~ use ~ the ~ formula~,
\\ A=PMT[\frac{(1+ \frac{r}{n})^{nt}-1}{\frac{r}{n}}] 
\\ Substitute~ the ~given ~data
\\A=250[\frac{(1+ \frac{0.115}{5})^{4(35)}-1}{\frac{0.115}{4}}]
\\ Simplify,
\\A=250[\frac{(1+ \frac{0.115}{5})^{105}-1}{\frac{0.115}{4}}]
\\A=250[\frac{(1+ 0.02875)^{105}-1}{0.0287}]
\\A=250[\frac{(1.02875)^{105}-1}{0.0287}]
\\A=250[\frac{19.61254997-1}{0.0287}]
\\A=250[\frac{18.61254997}{0.0287}]
\\A=250(647.39304)
\\A=161848.26061
\\ b) Yanmei ~has ~contributed ~\$250~ to ~an~ RRSP ~at ~the ~end ~of ~each ~3-month ~\\period ~for ~the~ past ~35 ~years.
\\ \therefore ~principle ~amount ~invested~ over ~35 ~year =[(35)(4)(250)]=\$35000
\\ \$35000~ of ~the~ investment ~will ~be~ interest~ earned ~over ~the~ 35 ~years.
\\~ and~
\\ Interest ~ earned=161848.26061-35000=126848.26061"